OpenStudy (anonymous):

-x-9y=39 Find the value of y when x=15

OpenStudy (anonymous):

y=-6

OpenStudy (anonymous):

You should be able to do these by now they are very simple just plug in the x value so your equation is -(15)-9y=39

OpenStudy (anonymous):

Put x = 15 -x -9y = 39 will become -15 -39 = 9y -54 = 9y y = -54/9